OpenStudy (anonymous):

Solve the system of equations by the Gauss-Jordan method. 1 1 1 | -8 1 2 2 | -17 1 3 2 | -22

OpenStudy (anonymous):

R2=R2-R1 R3=R3-R1 R3=R3-2R2 R3=-R3 R1=R1-R2 R2=R2-R3 Now you should be in reduced echelon form with solutions: x1=1,x2=-5,x3=-4